A sample of helium gas effuses at a rate of 12m/sec. Another gas effuses at a rate of 6.0m/sec. What is the molar mass of the unknown gas?
1
Expert's answer
2020-04-10T13:37:40-0400

R2/R1=√M1/M2

M1/(R2/R1)^2 = 4/(6/12)^2 =16 .

CH4